Rounding in the calculation of monthly interest rates is discouraged. Such rounding can lead to answers different from those presented here. For long-term loans, the differences may be pronounced.

You can get a car loan with a term of three years at an APR of 3%. If you can afford a monthly payment of $200, how much can you borrow? (Round your answer to the nearest cent.)
Answer #1

Amount to be borrowed=Monthly payment/(rate/12)*(1-1/(1+rate/12)^(12*3))=200/(3%/12)*(1-1/(1+3%/12)^(12*3))=6877.293020